OpenStudy (anonymous):

Explain why \[a_{n}\] monotically decreasing and lower bounded . Determine the limit \[lim_{n} a_{n}\] \[a_{n}=\frac{5n+1}{7n-2}\]

terenzreignz (terenzreignz):

I suppose we can consider the function \[\Large f(x) = \frac{5x+1}{7x-2}\]

terenzreignz (terenzreignz):

And find its derivative... \[\Large f'(x) = \frac{5(7x-2)-7(5x+1)}{(7x-2)^2}=-\frac{17}{(7x-2)^2}\] Which is negative for all values of x... (except \(\large\frac27\))

terenzreignz (terenzreignz):

There you have it, a negative derivative...

OpenStudy (anonymous):

did we showed in this case that is monotocially decreasing and lower bound ed?

terenzreignz (terenzreignz):

Well, its continuous counterpart is (strictly) decreasing for all positive integers...so its discrete case (the sequence) must be monotonically decreasing

OpenStudy (anonymous):

ok thx ;)